In developing empirically a cost function from observed data on a complex chemical experiment, an analyst employed normal error regression model (Yi = eta0 +eta1Xi+ varepsiloni ). eta0 was interpreted here as the cost of setting up the experiment. The analyst hypothesized that this cost should be $7.5 thousand and wished to test the hypothesis by means of a general linear test.

a) Indicate the alternative conclusions for the test

b) Specify the full and reduced models

c) Can you tell what the quantity dfR - dfF in test statistic SSE(R) - SSE(F)SSE(F) dfF   will equal in the analyst's test? Explain.
Answer #1

(a) The alternative conclusion for the test is the cost is should not be $7.5 thousand.

Ha: eta1 not equal to $7.5 thousand

(b) The full model or unrestricted model is the model thought to be most appropriate for the data. For simple linear regression the full model is:

Yi = eta0 +eta1Xi+ varepsiloni , where  Yi = is dependent variable  eta0 and eta1  are parameter estimates Xi is independent variable and

varepsiloni are errors in estimation

  • The reduced or restricted model is the model described by the null hypothesis Ho.

For simple linear regression , a common null hypothesis is H0:eta1 =0. or  H0: Yi = eta0 + varepsiloni.

In this case, the reduced model is taken when the null hypothesis is true.

The reduced model is given by:  Yi = eta0 + varepsiloni. where  eta0 is intercept of the model.

The reduced model suggests that the dependent or response variable Yi is a function of sum of overall mean , eta0 and error term varepsiloni  .

(c) Suppose we have 'n' observation for dependent variable X.

Then dfR is degree of freedom of reduced model (i.e. When the null hypothesis Ho:eta1 =0 is true )

and dfF is degree of freedom of full model (i.e. When the null hypothesis Ho:eta1 =0 is false )

Here,  dfR= (n-1), since in reduced model we are estimating on one parameter eta0. So degree of freedom = n- no of estimates = n-1 in this case.

dfF= (n-2), since in full model we are estimating on two parameters eta0 and eta1. So degree of freedom = n- no of estimates = n-2 in this case.

Hence, the quantity dfR - dfF =(n-1)-(n-2)= 1